# When do Neural Networks Learn World Models?

### v2

@@ 6 Limitations and Future Work
This work is an initial step towards formally understanding world model learning, and we can see many exciting future directions. (i) Neural networks in practice learn hierarchical representations (Bengio et al., 2013), which may require a more structured modeling of the representation $\Phi$. (ii) We do not assume the latent variables to have any specific structure (e.g., causality between variables); combining our analysis and the results in causal reasoning (Schölkopf et al., 2021; Richens & Everitt, 2024) would be interesting. (iii) We primarily focus on the low-complexity bias of neural networks; considering other implicit bias (Allen-Zhu & Li, 2023; Zhang et al., 2024) and more fine-grained complexity measures are also important directions for future work.
@@
This paper presents work whose goal is to advance our understanding of the inner workings of neural networks and LLMs, and our main results are of theoretical nature. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Related Work
To address this impossibility, subsequent research has explored various structural assumptions on latent variables, such as conditional independence between the latent variables and an observable auxiliary variable (Hyvarinen et al., 2019; Khemakhem et al., 2020; Lee et al., 2021), distributional constraints (Zimmermann et al., 2021; Wei et al., 2021), and causal interventions (Von Kügelgen et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023). Some studies have also linked these assumptions to contrastive learning (Hyvarinen et al., 2019; Tosh et al., 2021; Zimmermann et al., 2021). However, incorporating such structural assumptions often leads to complex and less scalable training paradigms compared to the pre-training framework of modern LLMs (i.e., next-token prediction). As a result, these studies do not directly address the central question of our work: Can the ongoing paradigm of LLMs learn world models? Meanwhile, the fact that LLMs already acquire non-trivial representations (Li et al., 2021; Bricken et al., 2023; Marks & Tegmark, 2023; Gurnee & Tegmark, 2024) suggests that they must leverage some form of implicit bias rather than explicit structural assumptions on input data, which motivates our study.
@@
A growing body of research introduces complexity measures tailored for neural networks trained via (stochastic) gradient descent (Bartlett et al., 2017; Jacot et al., 2018; Zhou et al., 2019; Lotfi et al., 2022; Chatterjee & Sudijono, 2024), often leveraging them for generalization analysis. However, these measures inherently depend on the neural network parameterization, rather than capturing function complexity independently of specific parameterizations. Some studies propose alternative complexity metrics inspired by Kolmogorov complexity in the machine learning context (Xu et al., 2020; Liu et al., 2023), but these metrics are typically problem-specific and do not enable a direct complexity analysis in the function space, unlike the approach we take in this work.
Appendix B Preliminaries on Boolean Function Analysis
@@ Appendix C Proofs
  • In Section C.1, we introduce some definitions and technical lemmas.
  • In Section C.2, we provide the proof of Theorem 4.1.
  • In Section C.3, we provide the proof of Theorem 4.3.
  • In Section C.4, we provide the proof of Theorem 4.4.
  • In Section C.5, we provide the proof of Corollary 4.7.
  • In Section C.6, we provide the proof of Theorem 4.8.
  • In Section C.7, we provide the proof of Theorem 4.9.
  • In Section C.8, we provide the proof of Theorem 4.12.
@@ C.1 Technical Lemmas
Proof of Lemma C.2.
We can prove Lemma C.2 by contradiction: assume that it does not hold, i.e., for some $S$, there exists ${\bm{x}}\in{\mathcal{X}}$ such that $\prod_{j\in[d+1]}(x_{i_{j}}+b_{j})\neq 0$, then we must have $x_{i_{j}}=b_{j},\forall j\in[d+1]$. Since $b_{1},\ldots,b_{d+1}$ are arbitrary, by applying this argument to every $(b_{1},\ldots,b_{d+1})\in\{-1,1\}^{d+1}$ we can find $2^{d+1}$ elements in ${\mathcal{X}}$ that differ in at least one coordinate in $S$, which yields $|{\mathcal{X}}|\geq 2^{d+1}$. On the other hand, recall that we assume $\mathrm{supp}({\mathbf{z}})={\mathcal{Z}}$. Due to the invertibility of $f$, we have $|{\mathcal{X}}|=|{\mathcal{Z}}|=2^{d}$, which contradicts $|{\mathcal{X}}|\geq 2^{d+1}$. Hence, the initial assumption is false and Lemma C.2 holds. ∎
Applying Lemma C.2, we have that for every degree-$d+1$ subset $S=\{i_{1},\ldots,i_{d+1}\}\subseteq[m]$,
@@
Remark C.3.
Without the latent structure, a trivial upper bound of the min-degree solution of any task $h$ is $\mathrm{deg}(\mathcal{H}_{\mathrm{min}}(h))\leq m$. Hence, Lemma C.1 is an example of how the min-degree bias can exploit the latent structure of data by favoring low-degree solutions with degree independent of the data dimension $m$.
@@
We first prove that for every $S$, $T$ is unique, by contradiction. Suppose that for some $S$, $T$ is not unique, i.e., there exists $T^{\prime}\neq T\subset[n]$ such that $T^{\prime}$ satisfies the condition. Then, by Definition C.4, changing the values of the coordinates $x_{i},i\in S$ can change only the values of $f_{j}({\bm{x}}),j\in T\cap T^{\prime}$ but not the values of other $f_{j}({\bm{x}}),j\in[n]\setminus(T\cup T^{\prime})$. This results in |$|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{n-|S|}\cdot 2^{|T\cap T^{\prime}|}=2^{n-k+|T\cap T^{\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and $T$ is unique. This allows us to define a mapping $\phi:S\mapsto T$.
Let ${\mathcal{S}}_{k}=\{S\subset[n]\mid|S|=k\}$ be the set of all subsets of $[n]$ with cardinality $k$. We then prove that $\phi:S\mapsto T$ is bijective on ${\mathcal{S}}_{k}$. To this end, it suffices to show that every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$. This can be similarly proved by contradiction as in proving the uniqueness of $T$: if it is false, then there exists a subset $S^{\prime\prime}=S\cup S^{\prime}$ and $T$ such that $\mathrm{Inf}_{i}(f_{j})=0,\,\forall i\in S^{\prime\prime},\,j\in[n]\setminus T$. Then, by Definition C.4, changing the values of the coordinates $x_{i},i\in S^{\prime\prime}$ can change only the values of $f_{j}({\bm{x}}),j\in T$ but not the values of other $f_{j}({\bm{x}}),j\notin T$. This results in |$|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{|T|}\cdot 2^{n-|S^{\prime\prime}|}=2^{n+k-|S^{\prime\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$.
@@
Lemma C.8 (Abbe et al. (2023), Theorem 5.1).
@@
Meanwhile, Definition 3.4 gives
@@
By Lemma C.1, we can upper-bound the degree of each $\Phi_{j}^{*},j\in[d]$ by
@@
Applying Lemma C.1, we have $\mathrm{deg}(\Phi)_{j}\leq d,\forall j\in[d]$. This gives
@@
By Lemma C.7, we have ${\mathcal{F}}^{d}\circ T^{-1}={\mathcal{F}}^{d}$. This gives
@@
for $g\circ\psi^{-1}\in{\mathcal{H}}(h)$. By Lemma C.1, we have $\mathrm{deg}(\mathcal{H}_{\mathrm{min}}(h))\leq d$ for every $h:{\mathcal{X}}\to{\mathbb{R}}$. Thus, for $\mathrm{deg}(\mathcal{H}_{\mathrm{min}}(h))>\mathrm{deg}(g)$ to hold, we must have $\mathrm{deg}(h\circ\psi)=\mathrm{deg}(g)\leq d-1$. By Definition 4.6, this gives $h\in{\mathcal{F}}^{d}_{d-1}\circ\psi^{-1}$, completing the proof. ∎
C.6 Proof of Theorem 4.8
@@
Proof of Lemma C.10.
For every ${\mathcal{G}}\subseteq{\mathcal{F}}^{d}$, let ${\mathcal{G}}\circ T=\{h^{\prime}\circ T\mid h^{\prime}\in{\mathcal{G}}\}$. By Lemma C.7, we know that the mapping $h^{\prime}\mapsto h^{\prime}\circ T$ is bijective on ${\mathcal{F}}^{d}$. We thus have $|{\mathcal{F}}^{d}_{k}\circ T|=|{\mathcal{F}}^{d}_{k}|$ for every $k$. For ${\mathcal{F}}^{d}_{k}\circ T$, there are two possibilities:
@@
Remark C.11.
A limitation of Theorem 4.8 is that we requires a non-zero probability of explicitly sampling degree-$1$ tasks (i.e., $p_{1}>0$), in which latent variables are essentially observed as task outputs. If $p_{1}=0$, we can still prove equation (61) by applying Lemma C.10; in other words, we can still prove that every representation $\Phi^{*}$ that learns the world model up to negations and permutations is a minimizer of the optimization problem (10). However, such minimizers may not be unique, because Lemma C.10 only proves the equivalence between the equality and $\mathrm{deg}(T_{i})=1,\forall i\in[d]$ when $k=1$ but not $11<k\leq d-$ 1. In fact, we can construct hard examples showing that in some cases, there indeed exists other $\Phi$ that minimizes $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime}\circ T)$ for every $k\in[d]\setminus\{1\}$.
Example C.12.
@@
One can easily verify that for every $k\in\{2,3\}$, every parity function $\chi_{S}(T({\bm{z}}))=\prod_{i\in S}T_{i}({\bm{z}})$ with $|S|\leq k$ satisfy $\mathrm{deg}(\chi_{S}(T({\bm{z}})))\leq k$. By the Fourier-Walsh transform, this amounts to ${\mathcal{F}}_{k}^{3}\circ T={\mathcal{F}}^{3}_{k}$ for $k=\{2,3\}$, which gives $\sum_{h^{\prime}\in{\mathcal{F}}^{3}_{k}}\mathrm{deg}(h^{\prime}\circ T)=\sum_{h^{\prime}\in{\mathcal{F}}^{3}_{k}}\mathrm{deg}(h^{\prime})$ by the proof of Lemma C.10. Thus, in this case we require $p_{1}>0$ to ensure that the representation $\Phi$ satisfying $\Phi({\bm{x}})=T^{-1}({\bm{z}})$ for every ${\bm{x}}\in{\mathcal{X}}$ and ${\bm{z}}=\psi^{-1}({\bm{x}})$ is not a minimizer of (10).
Nevertheless, we do believe that cases like Example C.12 are rare. This is because by the proof of Lemma C.10, such examples must construct a bijection $T$ such that ${\mathcal{F}}^{d}_{k}\circ T={\mathcal{F}}^{d}_{k}$ for every $k\in[d]\setminus\{1\}$, which is increasingly difficult when $d$ becomes large. For example, if we increase the dimension of ${\mathcal{Z}}$ from $3$ to $4$ in Example C.12 and keep $T_{1}$, $T_{2}$ and $T_{3}$ as is, it could be verified that there does not exist a $T_{4}:\{\-1,1\}^{3}\to\{-1,1\}$ satisfying that ${\mathcal{F}}^{4}_{k}\circ T={\mathcal{F}}^{4}_{k}$ for every $k\in\{2,3\}$. We believe that this intuition could be rigorously proved using e.g., Lemma C.6 or other techniques and leave it as future work.
C.7 Proof of Theorem 4.9
@@
Proof of Lemma C.13.
The main idea of the proof is similar to that of Lemma C.1 and Lemma C.2.
@@
Proof of (i). By Lemma C.13, for every $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$, we have $\mathrm{deg}(h^{*})\leq k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose r}\rceil$. The test MSE of any $h^{*}$ thus satisfies
@@
We then prove that $k\geq r+1$. To see this, recall that $d>r$ and one can directly verify
@@
Recall that $h\circ\psi$ is a parity function $\chi_{S}$ with $\mathrm{deg}(\chi_{S})=|S|=q\geq k$. Meanwhile, since $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$, by Definition 3.6 we have $\mathrm{deg}(h^{*})-\mathrm{deg}(g)\geq q-r$, i.e., $\mathrm{deg}(g)\leq\mathrm{deg}(h^{*})-q+r\leq r$, for every $g$ satisfying $g\circ\psi^{-1}=h^{*}$. This gives $\mathrm{deg}(h^{*}\circ\psi)\leq r$. Recall equation (19), we have
@@
Proof of (ii). Since $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and $\Phi^{*}$ learns the world model up to negations and permutations, we have $\mathrm{deg}(h\mid\Phi^{*})=\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and hence $\mathrm{deg}(g^{*})\leq r$. By Lemma C.8, $g^{*}$ is unique. We thus have $h=g^{*}\circ\Phi^{*}$, which gives the desired result. ∎
C.8 Proof of Theorem 4.12
@@
Proof of Lemma C.14.
By Defintion 4.11, we have $\mathrm{deg}(U(\chi_{S}))=\mathrm{deg}(\chi_{S})$ for every compatible $U$. Applying Lemma C.9, we further have $\mathrm{deg}(U(\chi_{S}))=\mathrm{deg}(\chi_{S})=\mathrm{deg}(U^{-1}(\chi_{S}))$. This gives

### v3

@@ 6 Limitations and Future Work
This work is an initial step towards formally understanding world model learning and the internal representations learned by neural networks, and we can see many exciting future directions. (i) Neural networks in practice learn hierarchical representations (Bengio et al., 2013), which may require a more structured modeling of the data generation model and the learned representations. (ii) We do not assume the latent variables to have any specific structure (e.g., causal relations between latent variables); extending our analysis to causal representation learning by combining our results and existing results in causal inference (Schölkopf et al., 2021; Richens & Everitt, 2024) is an interesting direction. (iii) We primarily focus on the low-complexity bias of neural networks; considering other implicit biases (Allen-Zhu & Li, 2023; Zhang et al., 2024) of neural networks and more fine-grained complexity measures are also promising directions for future work.
Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under STI 2030-Major Projects 2021ZD0200300, and in part by the National Key Research and Development Program of China No. 2024YDLN0006, and in part by the National Natural Science Foundation of China under Grant 62176133.
@@ Acknowledgements
This paper presents work whose goal is to advance our understanding of the inner workings of deep neural networks. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Related Work
To overcome this impossibility, subsequent research has explored various structural assumptions on latent variables, such as conditional independence between the latent variables and an observable auxiliary variable (Hyvärinen et al., 2019; Khemakhem et al., 2020; Lee et al., 2021), distributional constraints (Zimmermann et al., 2021; Wei et al., 2021), and causal interventions (Von Kügelgen et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023). Some studies have also linked these assumptions to contrastive learning (Hyvärinen et al., 2019; Tosh et al., 2021; Zimmermann et al., 2021). However, incorporating such structural assumptions often leads to complex and less scalable training paradigms compared to the pre-training framework of modern LLMs (i.e., next-token prediction). As a result, these studies do not directly address the central question of our work: Can the ongoing paradigm of LLMs learn world models? Meanwhile, the fact that LLMs already acquire non-trivial representations (Li et al., 2021; Bricken et al., 2023; Marks & Tegmark, 2023; Gurnee & Tegmark, 2024) suggests that they must leverage some form of implicit bias rather than explicit structural assumptions on input data, which motivates our study.
@@
A growing body of research introduces complexity measures tailored for neural networks trained via (stochastic) gradient descent (Bartlett et al., 2017; Jacot et al., 2018; Zhou et al., 2019; Lotfi et al., 2022; Chatterjee & Sudijono, 2024), often leveraging them for generalization analysis. However, these measures inherently depend on the neural network parameterization, rather than capturing function complexity independently of specific parameterizations. Some studies propose alternative complexity metrics inspired by Kolmogorov complexity in the machine learning context (Xu et al., 2020; Liu et al., 2023), but these metrics are typically problem-specific and do not enable a direct complexity analysis in the function space, unlike the approach we take in this work. See Section B for more discussion on some commonly-used complexity measures.
Appendix B Discussion on Complexity Measures
Besides the Boolean complexity measures considered by this work, there are also continuous approximations of Kolmogorov complexity, e.g., continuous complexity measures based on standard Fourier transform, the order of approximation polynomials, or compression (Xu et al., 2019; Pérez et al., 2019; Jiang et al., 2023; Teney et al., 2024). However, an important downside of these continuous complexity measures is that they do not apply to arbitrary continuous functions, neither constituting a basis of the continuous function space, which makes them less amenable to theoretical analysis.
For example, the linear function $f(x)=ax+b,x\in{\mathbb{R}}$ does not have a standard Fourier transform since it is not absolutely integrable. Even if we constrain it to some finite interval $[-L,L]$ and periodically extend it across the real axis, the resulting Fourier transform would involve frequencies up to infinity, suggesting infinite complexity if we use the highest frequency as the complexity measure. This is in contrast to the empirically observed simplicity bias of gradient descent, where linear models are often learned first (see e.g., Kalimeris et al. (2019)). The measure of polynomial order has a similar problem since not all continuous functions can be linearly represented by a set of polynomials. In fact, the degree of Boolean functions could be viewed as a “discrete” version of polynomial order, but it is more principled since the Boolean function space indeed has a polynomial basis given by the Fourier-Walsh transform. Finally, we found other compression-based approximations of Kolmogorov complexity hard to analyze theoretically since they rely on specific compression algorithms.
Appendix C Numerical Experiments
This section presents numerical experiments that corroborate our theoretical results. In all experiments, we train MLPs using stochastic gradient descent (SGD) and binary data from {±$\{\pm 1\}^{m}$, generated by the data generation model in Table 1 with $d=m=10$. A summarization of results is as follows:
  • Low-degree bias of neural networks: Figure 3 shows that MLPs indeed have a low-degree bias when solutions with different degrees exist for the same training distribution.
  • Impact of multi-task training on identifying data-generating variables: Figure 4 shows that the identification error of data-generating variables significantly decreases as more training tasks are drawn, while single-task learning does not lead to similar identification.
  • Identifying data-generating variables benefits OOD generalization: Figure 5 shows that in a Hamming sampling setting as in Theorem 4.9, MLPs with a representation $\Phi$ that identifies true data-generating variables achieve superior OOD generalization compared to MLPs without such a representation.
Please refer to the figures for more details and discussion.
It is worth noting that in all experiments, we use MLPs with quadratic activation $\sigma(x)=x^{2}$ to ensure the compatibility between the model and the Fourier-Walsh transform (i.e., the polynomial expression) of Boolean functions. Our preliminary experiments showed that the low-degree bias and the identification error are very sensitive to the activation function used, which is consistent with our basis compatibility analysis in Section 4.4. This also explains the phenomenon reported by Abbe et al. (2023) that MLPs with ReLU activation do not exhibit a “pure” low-degree bias for Boolean inputs.
Table 1: The data generation model ${\bm{x}}=\psi({\bm{z}})$ used in numerical experiments.
$x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $x_{6}$ $x_{7}$ $x_{8}$ $x_{9}$ $x_{10}$
$z_{1}$ $z_{1}z_{2}$ $z_{1}z_{2}z_{3}$ $z_{1}z_{2}z_{3}z_{4}$ $z_{1}z_{2}z_{3}z_{4}z_{5}$ $z_{2}z_{3}z_{4}z_{5}z_{6}$ $z_{3}z_{4}z_{5}z_{6}z_{7}$ $z_{4}z_{5}z_{6}z_{7}z_{8}$ $z_{5}z_{6}z_{7}z_{8}z_{9}$ $z_{6}z_{7}z_{8}z_{9}z_{10}$
(a) Task 1: $f({\bm{x}})=x_{1}x_{4}x_{5}$
(b) Task 2: $f({\bm{x}})=x_{1}x_{3}+0.5x_{1}x_{5}x_{6}$
Figure 3: Low-degree bias of neural networks: Training latents are uniformly sampled from the Hamming ball $B_{r}=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$; the out-of-distribution (OOD) test latents are uniformly sampled from {±$\{\pm 1\}^{d}\,(d=10)$. For small $r$, low-degree realizations of tasks (target functions) exist in the training distribution. A model that learns these low-degree realizations would have low in-distribution (ID) error yet high OOD error due to the loss of high-degree components in target functions. We observe that while SGD-trained MLPs learn ground-truth target functions when $r=10$ (denoted by “Full” in the legends), they indeed fit low-degree realizations for training data with $r=2$ and $r=3$, resulting in high OOD error. We separately train a three-layer MLP for each task and for each $r$.
(a) $2$-degree tasks
(b) $3$-degree tasks
(c) $4$-degree tasks
Figure 4: Impact of multi-task training on identifying data-generating variables: The identification error of data-generating variables decreases as the task number $n$ increases, measured by the MSE of the linear probe trained to fit true latent variables ${\mathbf{z}}$ given the learned representation $\Phi({\mathbf{x}})$. Each task is a randomly sampled parity function of ${\mathbf{z}}$ with degree not larger than $k$ for $k$-degree tasks. Compared to single-task training $(n=1)$, multi-task training with $n=64$ significantly reduces the identification error for all $k$-degree training task distributions with $k=2,3,4$, corroborating Theorem 4.3 and Theorem 4.8 in the main text. Both the representation $\Phi$ and each task-specific function $g_{i},\,i\in[n]$ are instantiated as a three-layer MLP.
(a) Task 1: $f({\bm{x}})=x_{1}x_{4}x_{5}$
(b) Task 2: $f({\bm{x}})=x_{1}x_{3}+0.5x_{1}x_{5}x_{6}$
Figure 5: Identifying data-generating variables benefits out-of-distribution generalization: For the same tasks considered in Figure 3 with $r=2$, SGD-trained MLPs on top of an oracle representation that identifies true data-generating variables (denoted by “w/ WM” in the legends) achieve out-of-distribution (OOD) generalization. By contrast, SGD-trained MLPs without the representation (denoted by “w/o WM” in the legends) fail. This corroborates Theorem 4.9 in the main text. We separately train an MLP for each task and for the setting with/without the oracle representation.
Appendix D Preliminaries on Boolean Function Analysis
@@ Appendix E Proofs
  • In Section E.1, we introduce some definitions and technical lemmas.
  • In Section E.2, we provide the proof of Theorem 4.1.
  • In Section E.3, we provide the proof of Theorem 4.3.
  • In Section E.4, we provide the proof of Theorem 4.4.
  • In Section E.5, we provide the proof of Corollary 4.7.
  • In Section E.6, we provide the proof of Theorem 4.8.
  • In Section E.7, we provide the proof of Theorem 4.9.
  • In Section E.8, we provide the proof of Theorem 4.12.
@@ E.1 Technical Lemmas
Proof of Lemma E.2.
We can prove Lemma E.2 by contradiction: assume that it does not hold, i.e., for some $S$, there exists ${\bm{x}}\in{\mathcal{X}}$ such that $\prod_{j\in[d+1]}(x_{i_{j}}+b_{j})\neq 0$, then we must have $x_{i_{j}}=b_{j},\forall j\in[d+1]$. Since $b_{1},\ldots,b_{d+1}$ are arbitrary, by applying this argument to every $(b_{1},\ldots,b_{d+1})\in\{-1,1\}^{d+1}$ we can find $2^{d+1}$ elements in ${\mathcal{X}}$ that differ in at least one coordinate in $S$, which yields $|{\mathcal{X}}|\geq 2^{d+1}$. On the other hand, recall that we assume $\mathrm{supp}({\mathbf{z}})={\mathcal{Z}}$. Due to the invertibility of $f$, we have $|{\mathcal{X}}|=|{\mathcal{Z}}|=2^{d}$, which contradicts $|{\mathcal{X}}|\geq 2^{d+1}$. Hence, the initial assumption is false and Lemma E.2 holds. ∎
Applying Lemma E.2, we have that for every degree-$d+1$ subset $S=\{i_{1},\ldots,i_{d+1}\}\subseteq[m]$,
@@
Remark E.3.
Without the latent structure, a trivial upper bound of the min-degree solution of any task $h$ is $\mathrm{deg}(\mathcal{H}_{\mathrm{min}}(h))\leq m$. Hence, Lemma E.1 is an example of how the min-degree bias can exploit the latent structure of data by favoring low-degree solutions with degree independent of the data dimension $m$.
@@
We first prove that for every $S$, $T$ is unique, by contradiction. Suppose that for some $S$, $T$ is not unique, i.e., there exists $T^{\prime}\neq T\subset[n]$ such that $T^{\prime}$ satisfies the condition. Then, by Definition E.4, changing the values of the coordinates $x_{i},i\in S$ can change only the values of $f_{j}({\bm{x}}),j\in T\cap T^{\prime}$ but not the values of other $f_{j}({\bm{x}}),j\in[n]\setminus(T\cup T^{\prime})$. This results in |$|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{n-|S|}\cdot 2^{|T\cap T^{\prime}|}=2^{n-k+|T\cap T^{\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and $T$ is unique. This allows us to define a mapping $\phi:S\mapsto T$.
Let ${\mathcal{S}}_{k}=\{S\subset[n]\mid|S|=k\}$ be the set of all subsets of $[n]$ with cardinality $k$. We then prove that $\phi:S\mapsto T$ is bijective on ${\mathcal{S}}_{k}$. To this end, it suffices to show that every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$. This can be similarly proved by contradiction as in proving the uniqueness of $T$: if it is false, then there exists a subset $S^{\prime\prime}=S\cup S^{\prime}$ and $T$ such that $\mathrm{Inf}_{i}(f_{j})=0,\,\forall i\in S^{\prime\prime},\,j\in[n]\setminus T$. Then, by Definition E.4, changing the values of the coordinates $x_{i},i\in S^{\prime\prime}$ can change only the values of $f_{j}({\bm{x}}),j\in T$ but not the values of other $f_{j}({\bm{x}}),j\notin T$. This results in |$|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{|T|}\cdot 2^{n-|S^{\prime\prime}|}=2^{n+k-|S^{\prime\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$.
@@
Lemma E.8 (Abbe et al. (2023), Theorem 5.1).
@@
On the other hand, Definition 3.4 gives
@@
By Lemma E.1, we can upper-bound the degree of each $\Phi_{j}^{*},j\in[d]$ by
@@
Applying Lemma E.1, we have $\mathrm{deg}(\Phi)_{j}\leq d,\forall j\in[d]$. This gives
@@
By Lemma E.7, we have ${\mathcal{F}}^{d}\circ T^{-1}={\mathcal{F}}^{d}$. This gives
@@
for $g\circ\psi^{-1}\in{\mathcal{H}}(h)$. By Lemma E.1, we have $\mathrm{deg}(\mathcal{H}_{\mathrm{min}}(h))\leq d$ for every $h:{\mathcal{X}}\to{\mathbb{R}}$. Thus, for $\mathrm{deg}(\mathcal{H}_{\mathrm{min}}(h))>\mathrm{deg}(g)$ to hold, we must have $\mathrm{deg}(h\circ\psi)=\mathrm{deg}(g)\leq d-1$. By Definition 4.6, this gives $h\in{\mathcal{F}}^{d}_{d-1}\circ\psi^{-1}$, completing the proof. ∎
E.6 Proof of Theorem 4.8
@@
Proof of Lemma E.10.
For every ${\mathcal{G}}\subseteq{\mathcal{F}}^{d}$, let ${\mathcal{G}}\circ T=\{h^{\prime}\circ T\mid h^{\prime}\in{\mathcal{G}}\}$. By Lemma E.7, we know that the mapping $h^{\prime}\mapsto h^{\prime}\circ T$ is bijective on ${\mathcal{F}}^{d}$. We thus have $|{\mathcal{F}}^{d}_{k}\circ T|=|{\mathcal{F}}^{d}_{k}|$ for every $k$. For ${\mathcal{F}}^{d}_{k}\circ T$, there are two possibilities:
@@
Remark E.11.
A limitation of Theorem 4.8 is that we requires a non-zero probability of explicitly sampling degree-$1$ tasks (i.e., $p_{1}>0$), in which latent variables are essentially observed as task outputs. If $p_{1}=0$, we can still prove equation (61) by applying Lemma E.10; in other words, we can still prove that every representation $\Phi^{*}$ that learns the world model up to negations and permutations is a minimizer of the optimization problem (10). However, such minimizers may not be unique, because Lemma E.10 only proves the equivalence between the equality and $\mathrm{deg}(T_{i})=1,\forall i\in[d]$ when $k=1$ but not $11<k\leq d-$ 1. In fact, we can construct hard examples showing that in some cases, there indeed exists other $\Phi$ that minimizes $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime}\circ T)$ for every $k\in[d]\setminus\{1\}$.
Example E.12.
@@
One can easily verify that for every $k\in\{2,3\}$, every parity function $\chi_{S}(T({\bm{z}}))=\prod_{i\in S}T_{i}({\bm{z}})$ with $|S|\leq k$ satisfy $\mathrm{deg}(\chi_{S}(T({\bm{z}})))\leq k$. By the Fourier-Walsh transform, this amounts to ${\mathcal{F}}_{k}^{3}\circ T={\mathcal{F}}^{3}_{k}$ for $k=\{2,3\}$, which gives $\sum_{h^{\prime}\in{\mathcal{F}}^{3}_{k}}\mathrm{deg}(h^{\prime}\circ T)=\sum_{h^{\prime}\in{\mathcal{F}}^{3}_{k}}\mathrm{deg}(h^{\prime})$ by the proof of Lemma E.10. Thus, in this case we require $p_{1}>0$ to ensure that the representation $\Phi$ satisfying $\Phi({\bm{x}})=T^{-1}({\bm{z}})$ for every ${\bm{x}}\in{\mathcal{X}}$ and ${\bm{z}}=\psi^{-1}({\bm{x}})$ is not a minimizer of (10).
Nevertheless, we do believe that cases like Example E.12 are rare. This is because by the proof of Lemma E.10, such examples must construct a bijection $T$ such that ${\mathcal{F}}^{d}_{k}\circ T={\mathcal{F}}^{d}_{k}$ for every $k\in[d]\setminus\{1\}$, which is increasingly difficult when $d$ becomes large. For example, if we increase the dimension of ${\mathcal{Z}}$ from $3$ to $4$ in Example E.12 and keep $T_{1}$, $T_{2}$ and $T_{3}$ as is, it could be verified that there does not exist a $T_{4}:\{\-1,1\}^{3}\to\{-1,1\}$ satisfying that ${\mathcal{F}}^{4}_{k}\circ T={\mathcal{F}}^{4}_{k}$ for every $k\in\{2,3\}$. We believe that this intuition could be rigorously proved using e.g., Lemma E.6 or other techniques and leave it as future work.
E.7 Proof of Theorem 4.9
@@
Proof of Lemma E.13.
The main idea of the proof is similar to that of Lemma E.1 and Lemma E.2.
@@
Proof of (i). By Lemma E.13, for every $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$, we have $\mathrm{deg}(h^{*})\leq k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose r}\rceil$. The test MSE of any $h^{*}$ thus satisfies
@@
We then prove that $k\geq r+1$. To see this, recall that $d>r$ and one can verify
@@
Since $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$, by Definition 3.6 we have $\mathrm{deg}(h^{*})-\mathrm{deg}(g)\geq q-r$ for every $g$ satisfying $g\circ\psi^{-1}=h^{*}$. This gives $\mathrm{deg}(g)\leq\mathrm{deg}(h^{*})-q+r\leq k-q+r$, i.e., $\mathrm{deg}(h^{*}\circ\psi)\leq k-q+r$. Recalling that $h\circ\psi$ is a parity function $\chi_{S}$ with $\mathrm{deg}(h\circ\psi)=|S|>k-q+r$ and applying equation (19), we have
@@
Proof of (ii). Since $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and $\Phi^{*}$ learns the world model up to negations and permutations, we have $\mathrm{deg}(h\mid\Phi^{*})=\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and hence $\mathrm{deg}(g^{*})\leq r$. By Lemma E.8, $g^{*}$ is unique. We thus have $h=g^{*}\circ\Phi^{*}$, which gives the desired result. ∎
E.8 Proof of Theorem 4.12
@@
Proof of Lemma E.14.
By Defintion 4.11, we have $\mathrm{deg}(U(\chi_{S}))=\mathrm{deg}(\chi_{S})$ for every compatible $U$. Applying Lemma E.9, we further have $\mathrm{deg}(U(\chi_{S}))=\mathrm{deg}(\chi_{S})=\mathrm{deg}(U^{-1}(\chi_{S}))$. This gives